OpenStudy (anonymous):

I need help on his equation! Solve for x: 1 over 3 (2x − 8) = 4

OpenStudy (anonymous):

2x-8+4*3 2x-8=12 2x=20 x=20/2 x=10

OpenStudy (anonymous):

10 is what I got at first, just wanted to make sure i was right.

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

Do you mind helping with other questions ?

OpenStudy (anonymous):

i can do that